OpenStudy (anonymous):

lim 1/(sqrt(1+x)-1) as x->0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ 1 }{ \sqrt{x+1}-1 } =\infty \]

OpenStudy (anonymous):

There is no other solution?

OpenStudy (anonymous):

Try Rationalizing .!

OpenStudy (anonymous):

I tried that

OpenStudy (anonymous):

Still infinity

OpenStudy (anonymous):

With L'Hopital I get 0

OpenStudy (anonymous):

But I need solution without it